INCOME STATEMENT BALANCE SHEET
  Sales $ 26,700   Assets $ 119,000      Debt $ 27,800   
  Costs 16,400   Equity 91,200   
  Taxable income $ 10,300       Total $ 119,000         Total $ 119,000   
  Taxes (32%) 3,296
      Net income $ 7,004

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $3,242.8 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $37,113. What is the external financing needed?
Answer #1

Growth rate in sales=(37113-26700)/26700=39%

Dividend payout ratio=Dividend/Net income

(3242.8/7004)=0.462992575

Sales 37113
Costs(16400*1.39) 22796
Taxable income $14317
Taxes(32%*$14317) $4581.44
Net income $9735.56
Less:Dividends($9735.56*0.462992575) 4507.492
Addition to Retained earnings $5228.068

Total Assets would be=$119,000*1.39=$165410

Total equity=$91200+Addition to retained earnings

=(91200+5228.068)=$96428.068

Total Assets=Total equity+Total debt

Hence external financing needed=$165410-(96428.068+27800)

=$41181.932